OpenStudy (anonymous):

simplify the expression (sqrt5-sqrt6)(sqrt5+sqrt6)

OpenStudy (anonymous):

\[(\sqrt{5}-\sqrt{6})(\sqrt{5}+\sqrt{6})\] \[5+\sqrt{30}-\sqrt{30}-6\] when you do a binomial expansion collect like terms and you have your answer

OpenStudy (anonymous):

Will do you need more explanation?

OpenStudy (anonymous):

(sqrt5-sqrt6)(sqrt5+sqrt6) It's in the form (a -b) ( a + b) = a^2 +b^2 -> 5 - 6 = -1